A gardener has 25 kilos of a mixture of fertilizer and weed killer. If 1 kilo of the mixture is replaced by weed killer, the result is a mixture that is 10% weed killer. What percent of the original mixture was weed killer?
Answer by ankor@dixie-net.com(22740)

A gardener has 25 kilos of a mixture of fertilizer and weed killer.
If 1 kilo of the mixture is replaced by weed killer, the result is a mixture that is 10% weed killer.
What percent of the original mixture was weed killer?
:
let x = the percent of weed killer in the original mixture (in decimal form)
x(25 - 1) + 1 = .10(25)
25x - x + 1 = 2.5
24x + 1 = 2.5
24x = 2.5 - 1
24x = 1.5
x = 1.5/24
x = .0625 = 6.25% is the original percent